|5/2 - x| = |2 - x/2|
Answer by stanbon(75887)

|5/2 - x| = |2 - x/2|
---
(5/2)-x = 2 -(x/2) or (5/2)-x = -(2-(x/2))
(x/2) = 1/2 or (5/2)-x = (x/2)-2
x = 1 or (3/2)x = 9/2
x = 1 or x = 3
---------------------
Checking:
Does x = 1?
|5/2 - 1| = |2- 1/2|
3/2 = 3/2
-------------------
Does x = 3
|5/2 - 3| = |2- (3/2)|
|-1/2| = |1/2|
1/2 = 1/2
=================
Cheers,
Stan H.